OpenStudy (anonymous):

Use common denominators to determine whether the ratios form a proportion. 6 : 9 and 9 : 12 A. Yes. They form a proportion. B. No. They do not form a proportion. is it B?

OpenStudy (anonymous):

@k_lynn @Data_LG2

OpenStudy (anonymous):

Yes, B is correct.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We can change them to denominators of 108. \(\dfrac{6}{9} \rightarrow \dfrac{6 \times 12}{9 \times 12} \rightarrow \dfrac{72}{108}\) \(\dfrac{9}{12} \rightarrow \dfrac{9 \times 9}{12 \times 9} \rightarrow \dfrac{81}{108}\) Since \(\dfrac{72}{108}\) is less than \(\dfrac{81}{108}\) they are not the same.

OpenStudy (anonymous):

Thank u

OpenStudy (anonymous):

I have a few more

OpenStudy (anonymous):

Use division to determine whether the ratios form a proportion. 1/4 6/24. and A. Yes. They form a proportion. B. No. They do not form a proportion.

OpenStudy (anonymous):

Let mesolve problem

OpenStudy (anonymous):

@iGreen. I don't know how to do this one please help

OpenStudy (anonymous):

1 question per post, close this and open a new one.

OpenStudy (anonymous):

Okay sorry